# CoTracker: It is Better to Track Together

### v2

@@ 3.2 Transformer architecture and virtual tracks
The transformer $\Psi$ interleaves attention operators that operate across the time and track dimensions, respectively. Factorising the attention [1] across time and tracks makes the model computationally tractable: the complexity is reduced from $O(N^{2}T^{2})$ to $O(N^{2}+T^{2})$. However, for very large values of $N$, this cost is still prohibitive, especially when tracking points densely in a video. We thus propose a new design, in which we introduce $K$ virtual tracks tokens, where $K\ll N$ is a hyperparameter. These are just added as additional tokens to the input grid with a fixed, learnable initialization and removed at the output. They participate in time attention as regular tracks, and in cross attention with all other tracks, reducing the cost to $O(NK+K^{2}+T^{2})$.
Starting with DETR [6], various computer vision transformers have used learnable queries. For instance, [20] uses them to decode dense outputs. Our virtual tracks are different, however: they mirror and shadow the computation of the prediction targets (i.e., the tracks), actively participating in the calculations to reduce the computational burden.
3.3 Windowed inference
@@
A key advantage of formulating tracking as iterated updates is that it can easily support a sliding window application to process long videos. Consider in particular a video $V$ of length $T^{\prime}>T$ longer than the maximum window length $T$ supported by the architecture. To track points throughout the entire video $V$, we split the video in $J=\left\lceil 2T^{\prime}/T-1\right\rceil$ windows of length $T$, with an overlap of $T/2$ frames. [^1] We assume that $T$ is even. The last window is shorter if $T/2$ does not divide $T^{\prime}$.
@@
Let the superscript $(m,j)$ denote the $m$-th application of the transformer to the $j$-th window. We thus have a $M\times J$ grid of quantities $(\hat{P}^{(m,j)},~{}\hat{v}^{(m,j)},~{}Q^{(m,j)})$, spanning transformer iterations and windows. Quantities $m=0$ and $j=1$ are initialized as for the single window case. Then, the transformer is applied $M$ times to obtain the estimate $(M,1)$. The latter is used to initialize estimate $(0,2)$ by broadcasting the known quantities as before. Specifically, the first $T/2$ components of $\hat{P}^{(0,2)}$ are copies of the last $T/2$ components of $\hat{P}^{(M,1)}$; the last $T/2$ components of $\hat{P}^{(0,2)}$ are instead copies of the last time $t=T/2-1$ from $\hat{P}^{(M,1)}$. The same update rule is used for $\hat{v}^{(0,2)}$, while $Q^{(0,j)}$ is always initialized with initial track features $Q$. This process is repeated until estimate $(M,J)$ is obtained.
3.4 Unrolled window training
We found it important to learn the windowed transformer in an unrolled fashion in order to learn to handle semi-overlapping windows. The primary loss is for track regression, summed over iterated transformer applications and windows:
@@ 3.4 Unrolled window training
where $\gamma=0.8$ discounts early transformer updates. Here $P^{(j)}$ contains the ground-truth trajectories restricted to window $j$ (trajectories which start in the middle of the window are padded backwards). The second loss is the cross entropy of the visibility flags $\mathcal{L}_{2}(\hat{v},v)=\sum_{j=1}^{J}\operatorname{CE}(\hat{v}^{(M,j)},v^{(j)})$. While only a moderate number of windows are used in the loss during training due to the computational cost, at test time we can unroll the windowed transformer applications arbitrarily, thus in principle handling any video length.
@@
A technical difficulty is that the number of tracks $N$ is not fixed, but varies from window to window as new points can be added to the tracker at any time. While conceptually this is handled by simply adding more tokens when needed, in practice changing the number of tokens makes batching data for training difficult. Instead, we allocate enough tokens to support all points that must be tracked, regardless of which window they are added in, and use masking to ignore tokens that are not yet used.
3.5 Support points
CoTracker can take advantage of tracking several points jointly. It is typical for applications to have several points to track, but in some cases, one might be interested in tracking a few or even a single point. In fact, single-point tracking is a case we consider in the evaluation for fairness (Sec. 1).
In practice, we found it beneficial to track additional ‘support points’ which are not explicitly requested by the user, particularly when the goal is to track only a few of them. Moreover, we have found that different configurations of such support points can lead to small differences in performance. We experiment with various configuration types, visualized in Fig. 4. With the “global” strategy, the support points form a regular grid across the whole image. With the “local” strategy, the grid of points is centred around the point we wish to track, thus allowing the model to focus on a neighbourhood of it. We also test using the SIFT detector [26] to select support points. Note that these patterns are only considered at inference time and are used to improve the accuracy of the tracker for the target points.
4 Experiments
We thoroughly evaluate CoTracker on standard real and synthetic tracking benchmarks to assess its performance and generalization properties in difficult conditions and against numerous state-of-the-art trackers.
Method Training data DAVIS First DAVIS Strided Dynamic Replica
AJ $\uparrow$ $\delta_{\text{avg}}^{\text{vis}}$uparrow OA $\uparrow$ AJ $\uparrow$ $\delta_{\text{avg}}^{\text{vis}}$uparrow OA $\uparrow$ $\delta_{\text{avg}}$uparrow $\delta_{\text{avg}}^{\text{vis}}$uparrow $\delta_{\text{avg}}^{\text{occ}}$uparrow
TAP-Net [10] TAP-Vid-Kubric 33.0 48.6 78.8 38.4 53.1 82.3 45.5 53.3 20.0
OmniMotion [46] 51.7 67.5 85.3
PIPs [15] FlyingThings++ 42.2 64.8 77.7 52.4 70.0 83.6 41.0 47.1 21.0
MFT [31] Kubric + others 47.3 66.8 77.8 56.1 70.8 86.9
PIPs++ [52] PointOdyssey 69.1 73.7 55.5 64.0 28.5
TAPIR [11] TAP-Vid-Kubric 56.2 70.0 86.5 61.3 73.6 88.8 56.8 66.1 27.2
CoTracker (Ours) TAP-Vid-Kubric 62.2 75.7 89.3 65.9 79.4 89.9 61.6 68.9 37.6
Table 1: State of the art comparison. We compare CoTracker to the best trackers available on TAP-Vid-DAVIS utilizing both their “queried first” and “queried strided” evaluation protocols as well as on Dynamic Replica. During evaluation on DAVIS, TAPIR and PIPs++ have access to all the video frames at once, while we process videos in an online manner, using a causal sliding window.
@@ 4 Experiments
Table 2: PointOdyssey. All methods are trained on PointOdyssey. ${}^{\dagger}$ results for [52] obtained using released code and model.
Attention Points DAVIS First
time joint target
support
AJ $\uparrow$ $\delta_{\text{avg}}$uparrow OA $\uparrow$
(a) single
58.3 71.5 86.4
(b) single
41.1 62.9 75.1
(c) single
glob. $9\!\times\!9$
56.8 71.2 85.8
(d) single
SIFT $8\!\times\!8$
57.4 72.1 85.8
(e) single
loc. $10\!\times\!10$
60.4 75.2 87.5
(f) single
glob. $5\!\times\!5$ +loc. $8\!\times\!8$
62.2 75.7 89.3
(g) all
57.6 73.2 86.1
(h) all
glob. $5\!\times\!5$
60.5 75.8 87.6
(i) all
SIFT $8\!\times\!8$
60.7 75.7 88.1
Table 3: Importance of joint tracking. We compare using time and cross-track attention, tracking single or multiple target points, and using additional support points.
@@
Table 5: The virtual tracks allow CoTracker to scale. We report the maximum number of tracks that can fit on a 80 GB GPU.
Table 6: Single point vs. joint tracking. We qualitatively compare tracking individual points (top row) to tracking all points together (bottom). Background points are colored cyan while foreground points are magenta. We find that single-point tracking leads to background points following the object motion, as the model cannot reason well about the global context of the video (cyan points in later frames). Tracking points together results in better tracks for foreground and background points.
Num. virt. tracks DAVIS First Max. num. tracks Time [s]
@@
0 61.6 75.6 88.3 $97\!\times\!97$ 207.3
32 60.2 74.4 88.5 $263\!\times\!263$ 26.8
64 62.2 75.7 89.3 $263\!\times\!263$ 27.9
128 60.9 74.8 88.4 $263\!\times\!263$ 30.1

### v3

@@ 3.2 Transformer architecture and proxy tokens
The transformer $\Psi$ interleaves attention layers operating across the time and track dimensions, respectively. Factorising the attention [2] across time and tracks makes the model computationally tractable: the complexity is reduced from $O(N^{2}T^{2})$ to $O(N^{2}+T^{2})$. However, for very large values of $N$, this cost is still prohibitive. We thus propose a new design, in which we introduce $K$ proxy tracks, where $K\ll N$ is a hyper-parameter. Proxy tracks are learned fixed tokens that we concatenate to the list of ‘regular’ tracks at the input to the transformer and discard at the output.
For time attention, ‘regular’ and proxy tracks are processed identically. For track attention, however, regular tracks cross-attend the proxies, but not each other, reducing the cost to $O(NK+K^{2}+T^{2})$. We discard the proxy tracks at the output of the transformer.
Proxies are often used to accelerate, e.g., large graph neural networks; here, proxies are learnable query tokens. In computer vision, DETR [9] makes use of learnable queries in a transformer and methods like [30] use them to decode dense outputs. Our proxy tokens are different: They mirror and shadow the computation of the regular tracks, similarly to registers [17].
3.3 Windowed inference and unrolled training
An advantage of formulating tracking as the progressive refinement of a window of tracks is the ability to process arbitrarily long videos: It suffices to initialize the next window using partially overlapping tracks from the previous one.
@@ 3.3 Windowed inference and unrolled training
Consider, in particular, a video $V$ of length $T^{\prime}>T$ longer than the maximum window length $T$ supported by the architecture. To track points throughout the entire video $V$, we split the video in $J=\left\lceil 2T^{\prime}/T-1\right\rceil$ windows of length $T$, with an overlap of $T/2$ frames. [^1] We assume that $T$ is even. The last window is shorter if $T/2$ does not divide $T^{\prime}$.
@@
Let the superscript $(m,j)$ denote the $m$-th application of the transformer to the $j$-th window. We thus have a $M\times J$ grid of quantities $(\hat{P}^{(m,j)},~{}\hat{v}^{(m,j)},~{}Q^{(m,j)})$, spanning transformer iterations and windows. For $m=0$ and $j=1$, these are initialized as for the single window case. Then, the transformer is applied $M$ times to obtain the estimate $(M,1)$. The latter is used to initialize estimate $(0,2)$ via broadcasting. Specifically, the first $T/2$ components of $\hat{P}^{(0,2)}$ are copies of the last $T/2$ components of $\hat{P}^{(M,1)}$; the last $T/2$ components of $\hat{P}^{(0,2)}$ are instead copies of the last time $t=T/2-1$ from $\hat{P}^{(M,1)}$. The same update rule is used for $\hat{v}^{(0,2)}$, while $Q^{(0,j)}$ is always initialized with the features $Q$ of the track’s staring location. This process is repeated until estimate $(M,J)$ is obtained.
The windowed transformer essentially operates as a recurrent network, so we train it in an unrolled fashion. Specifically, we optimize the track prediction error summed over iterated transformer applications and windows
@@
where $\gamma=0.8$ discounts early transformer updates. Here, $P^{(j)}$ contains the ground-truth trajectories restricted to window $j$ (trajectories which start in the middle of the window are padded backwards). The second loss is the cross entropy of the visibility flags $\mathcal{L}_{2}(\hat{v},v)=\sum_{j=1}^{J}\operatorname{CE}(\hat{v}^{(M,j)},v^{(j)})$. While only a moderate number of windows are used in the loss during training due to the computational cost, at test time, we can unroll the windowed transformer applications arbitrarily, thus, in principle, handling any video length. In combination with joint tracking, unrolling allows the model to track points through occlusions of duration longer than one sliding window and to track points even outside of the camera view.
3.4 Support points
CoTracker can take advantage of tracking several points jointly. It is typical for applications to have several points to track, but in some cases, one might be interested in tracking a few or even a single point at inference time.
In these cases, we found it beneficial to track additional support points which are not explicitly requested by the user. Moreover, we have found that different configurations of such support points can lead to small differences in performance. We experiment with various configuration types, visualized in Fig. 5. With the “global” strategy, the support points form a regular grid across the whole image. With the “local” strategy, the grid of points is centred around the point we wish to track, thus allowing the model to focus on a neighbourhood of it. Note that these patterns are only considered at inference time and are used to improve the tracker’s accuracy for the target points by incorporating context.
4 Experiments
Figure 5: Tracked and support points. We visualize different configurations of points along with their tracks. Points represent the start of each track, which clearly illustrates grids of points in (b) and (c). (a) A single target point in a sequence from TAP-Vid-DAVIS. (b) A global grid of support points. (c) A local grid of support points. (d) Local and global support points. In Tab. 7 we evaluate CoTracker by combining different configurations of support points, as described in Sec. 3.4.
We evaluate CoTracker on several standard real and synthetic tracking benchmarks. After discussing implementation details, datasets and benchmarks, we compare CoTracker to the state of the art (Sec. 4.1) and ablate our design, demonstrating the importance of joint tracking and other innovations (Sec. 4.2)
Datasets and benchmarks.
We train the model on TAP-Vid-Kubric [18], as in [18, 19]. It consists of sequences of 24 frames showing 3D rigid objects falling to the ground under gravity and bouncing, generated using the Kubric engine [23]. Point tracks are selected randomly, primarily on objects and some on the background. Objects occlude each other and the background as they move.
For evaluation, we consider TAP-Vid-DAVIS [18], containing 30 real sequences of about 100 frames, and TAP-Vid-RGB-Stacking [18] with 50 synthetic sequences with objects being moved by robots of 200–300 frames each. Points are queried on random objects at random times and evaluation assesses both predictions of positions and visibility flags. This uses two evaluation protocols. In the “first” protocol, each point is queried only once in the video, at the first frame where it becomes visible. The tracker is expected to operate causally, predicting positions only for future frames. In the “strided” protocol, points are queried every five frames, and tracking is bidirectional. Given that most trackers (ours, PIPs, PIPs++) are causal, we run them twice, on the video and its reverse. We assess tracking accuracy using the TAP-Vid metrics [18]: Occlusion Accuracy (OA; accuracy of occlusion prediction treated as binary classification), $\delta_{\text{avg}}^{\text{vis}}$ (fraction of visible points tracked within 1, 2, 4, 8 and 16 pixels, averaged over thresholds), and Average Jaccard (AJ, measuring jointly geometric and occlusion prediction accuracy). Following [18], the $\delta_{\text{avg}}^{\text{vis}}$ and AJ tracking thresholds are applied after virtually resizing images to $256\times 256$ pixels. Note that TAP-Vid benchmarks evaluate only visible points.
We also evaluate on PointOdyssey [69], a more recent synthetic benchmark dataset for long-term tracking. It contains 100 sequences that are several thousand frames long, with objects and characters moving around the scene. Scenes are much more realistic than TAP-Vid-Kubric. We train and evaluate CoTracker on PointOdyssey and report $\delta_{\text{avg}}^{\text{vis}}$ and $\delta_{\text{avg}}^{\text{occ}}$, $\delta_{\text{avg}}$. The last two are the same as $\delta_{\text{avg}}^{\text{vis}}$, but for occluded and all points, respectively; they can be computed because the dataset is synthetic so the ground-truth positions of invisible points are known. We also report their Survival rate, which is the average fraction of video frames until the tracker fails (detected when the tracking error exceeds 50 pixels).
Dynamic Replica [35] is a synthetic dataset for 3D reconstruction that contains long-term tracking annotations. It consists of 500 300-frame sequences of articulated models of people and animals. We evaluate models trained on Kubric and measure $\delta_{\text{avg}}^{\text{vis}}$ and $\delta_{\text{avg}}^{\text{occ}}$ on the “valid” split of this dataset.
@@ 4 Experiments
Here we provide important implementation details and refer to the sup. mat. for others. We train the model for 50,000 iterations on 6,000 TAP-Vid-Kubric sequences of $T^{\prime}=24$ frames using sliding window size $T=8$, which takes 40 hours. We use $32$ NVIDIA A100 80GB GPUs with a batch size of 1. Training tracks are sampled preferentially on objects. During training, we construct batches of $N=768$ tracks with visible points either in the first or middle frames of the sequence to train the tracker to handle both cases. In a similar manner, we also train a second version of CoTracker on the train split of PointOdyssey, using 128 tracks randomly sampled from sequences of length $T^{\prime}=56$.
A technical difficulty is that the number of tracks $N$ varies from window to window, as new points can be added to the tracker at any time. While conceptually, this is handled by simply adding more tokens when needed, in practice, changing the number of tokens makes batching data for training difficult. Instead, we allocate enough tokens for all the tracked points, regardless of which window they are added in, and use masking to ignore tokens that are not yet used.
The video resolution can also affect tracking performance. For evaluation on TAP-Vid, we follow their protocol and downsample videos to $256\times 256$, which ensures that no more information than expected is passed to the tracker. However, each tracker has a different native resolution at which it is trained ($384\times 512$ for PIPs and CoTracker; $512\times 896$ for PIPs++). We thus resize videos from $256\times 256$ to the native resolution of each tracker before running it to ensure fairness. For Dynamic Replica and PointOdyssey, we resize videos to the native tracker resolution, as there is no prescribed resolution.
Method Train DAVIS First DAVIS Strided RGB-S First
[…]
TAP-Net [18] K 33.0 48.6 78.8 38.4 53.1 82.3
OmniMotion [62] 52.8 66.9 87.1 51.7 67.5 85.3 69.5 82.4 90.3
PIPs [24] FT++ 42.2 64.8 77.7 52.4 70.0 83.6 59.1
MFT [45] K++ 47.3 66.8 77.8 56.1 70.8 86.9
PIPs++ [69] PO 69.1 73.7 77.8
TAPIR [19] K 56.2 70.0 86.5 61.3 73.6 88.8 60.3 74.1 85.5
CoTracker (Ours) K 62.2 75.7 89.3 65.9 79.4 89.9 71.6 83.3 89.6
Table 1: TAP-Vid benchmarks We compare CoTracker to the best trackers available on TAP-Vid benchmarks utilizing both their “first” and “strided” evaluation protocols. During evaluation, TAPIR and OmniMotion have access to all video frames at once, while we process videos in an online manner using a causal sliding window. This puts online trackers like ours at a disadvantage on long videos with slow motion, such as in RGB-S. To address this, we decrease the frame rate for RGB-S evaluation by keeping every 5th frame. Training data: (K) Kubric, (K++) Kubric+more, (FT) FlyingThings++, (PO) Point Odyssey.
@@
Table 2: PointOdyssey. All methods are trained on PointOdyssey. results for [69] obtained using released code and model.
Method Dynamic Replica
$\delta_{\text{avg}}$uparrow $\delta_{\text{avg}}^{\text{vis}}$uparrow $\delta_{\text{avg}}^{\text{occ}}$uparrow
TAP-Net [18] 45.5 53.3 20.0
PIPs [25] 41.0 47.1 21.0
PIPs++ [69] 55.5 64.0 28.5
TAPIR [19] 56.8 66.1 27.2
CoTracker (Ours) 61.6 68.9 37.6
Table 3: Dynamic Replica. The gap between CoTracker and other methods in tracking occluded points’ accuracy $\delta_{\text{avg}}^{\text{occ}}$ is higher than $\delta_{\text{avg}}^{\text{vis}}$.
4.1 Comparisons to the State of the Art
We compare CoTracker to state-of-the-art trackers on TAP-Vid-DAVIS and TAP-Vid-RGB-Stacking in Tab. 1. For fairness, we use the same configuration of support points identified in the ablations (see Sec. 4.2) across all benchmarks. We also evaluate CoTracker on PointOdyssey in Tab. 3 and on Dynamic Replica in Tab. 3 by tracking jointly all the target points. CoTracker improves most metrics by substantial margins on the TAP-Vid benchmarks while generalizing well from synthetic Kubric to real DAVIS. CoTracker also improves tracking accuracy $\delta_{\text{avg}}^{\text{vis}}$ in all cases. On PointOdyssey, despite using a window size of only 8 frames, CoTracker has a better survival rate than PIPs++, which uses a 128-frame sliding window. This shows the power of unrolled training, which trains the model to propagate information across a long series of sliding windows. The gap between CoTracker and other methods in Tab. 3 and Tab. 3 for $\delta_{\text{avg}}^{\text{occ}}$ is higher than for $\delta_{\text{avg}}^{\text{vis}}$, which shows that CoTracker excels in tracking occluded points thanks to jointly tracking groups of points.
4.2 Ablations
Tracking together is better.
In Tab. 5, we demonstrate the importance of tracking points jointly, which is the main motivation of CoTracker. This is achieved by removing cross-track attention from the tracker, thus ignoring the dependencies between tracks entirely. For fairness, rather than simply removing the six cross-track attention layers, we maintain the same model size and replace them with twelve time-attention layers, resulting in a comparable model size. On Dynamic Replica, joint tracking improves the accuracy of occluded points $\delta_{\text{avg}}^{\text{occ}}$ 28.8 $\rightarrow$ 37.6 (+30.6%) more than the accuracy of visible points $\delta_{\text{avg}}^{\text{vis}}$ 62.4 $\rightarrow$ 68.9 (+10.4%), showing the effectiveness of joint tracking in understanding the scene motion, including currently invisible points.
Mode: DAVIS First Dynamic Replica
AJ $\uparrow$ $\delta_{\text{avg}}$uparrow OA $\uparrow$ $\delta_{\text{avg}}$uparrow $\delta_{\text{avg}}^{\text{vis}}$uparrow $\delta_{\text{avg}}^{\text{occ}}$uparrow
no joint 55.6 70.1 83.0 54.4 62.4 28.8
joint 62.2 75.7 89.3 61.6 68.9 37.6
Table 4: Importance of joint tracking. We track a single target point with or without additional support points on DAVIS, and all target points separately or jointly on Dynamic Replica.
@@ 4.2 Ablations
Importance of unrolled training.
We design CoTracker for tracking in a windowed manner, and in Tab. 5, we thus assess the effect of unrolling windows during training (Sec. 3.3). Switching off unrolled training decreases performance by 18 AJ points. Hence, unrolled training helps to track over long periods of time, >10×>10$\times$> 10 × longer than the sequences used in training.
Effect of proxy tokens on scalability.
In Tab. 7, we assess the benefits of using proxy tokens during inference. For a fixed memory budget (80 GB), using proxy tokens instead of full self-attention allows us to track ×$\times 7.4$ more points than without them; in fact, we can track a whole 263 $\times$ 263 grid, which is quasi-dense for the input video resolution. Moreover, proxy tokens also reduce the time complexity of the model and make it ×$\times 7$ faster at inference time for the maximum number of tracked points. The number of proxy tokens does not affect scalability but affects performance, with the best results obtained using 64 proxy tokens. In short, proxy tokens allow to track almost one order of magnitude more tracks with higher accuracy than the naive self-attention.
Num. proxy tokens DAVIS First Max. num. tracks Time [s]
@@
0 61.6 75.6 88.3 $9.4$k 207.3
32 60.2 74.4 88.5 $69.2$k 26.8
64 62.2 75.7 89.3 $69.2$k 27.9
128 60.9 74.8 88.4 $69.2$k 30.1
Table 6: The proxy tokens allow CoTracker to scale. We train models with different numbers of proxy tokens and show that they help to decrease both memory complexity and inference time. We report the maximum number of tracks that can fit on a 80 GB GPU.
Support: DAVIS First
global local AJ $\uparrow$ $\delta_{\text{avg}}$uparrow OA $\uparrow$
55.6 70.1 83.0
56.8 71.2 85.8
60.4 75.4 87.3
62.2 75.7 89.3
Table 7: Support points for TAP-Vid benchmarks. We track single target points from DAVIS using additional support points for context, as described in Sec. 3.4. Adding points for context always helps, while additional global and local context works best.
Optimal support point configurations.
We compare the effect of choosing different configurations of support points from Fig. 5 in Tab. 7. In order to do so, we take a single benchmark point at a time and add additional support points (Sec. 3.4) to allow the model to perform joint tracking. This protocol, where a single benchmark point is tracked each time, is also important for fairness when compared to the state-of-the-art. This is because the benchmark points might be biased for objects, which is particularly evident in TAP-Vid-DAVIS (Fig. 5). Passing more than one such biased point at the same time to a joint tracker like ours could help it by revealing the outline of objects to the tracker. By considering one benchmark point at a time, we ensure that no such ground-truth information can leak into it. We use this scheme for the TAP-Vid benchmarks, which explicitly prohibits tracking more than one ground-truth point at a time.
Adding any configurations of support points helps, but the local configuration helps much more than the global, presumably because the model uses dependencies with other points located on the same object. The combination of global and local context works best, likely because the model can track both the camera motion and the object motion in this case.
Figure 7: Qualitative Results. CoTracker can track points outside of the camera view. We visualize point displacements on a dense regular grid using an example from DAVIS [46]. In this example, the camera moves towards a group of people, while points in the background stay behind the camera view.
4.3 Limitations
Despite its high performance, CoTracker still occasionally makes tracking mistakes that a human would not make. Since CoTracker is trained on purely synthetic data, it sometimes does not generalize well to complex visual scenes with reflections and shadows. For example, CoTracker tends to track shadows together with the objects that cast them. Depending on the application this may be a desirable property (e.g. video editing) or it may not be (e.g. motion analysis).
5 Conclusions
We have presented CoTracker, a transformer-based point tracker that tracks several points jointly, accounting for their dependencies. CoTracker is state-of-the-art on the standard tracking benchmarks, often by a substantial margin, can track through occlusions and when points leave the field of view, even for hundreds of frames, and can track a very large number of points simultaneously. The transformer architecture is flexible and memory efficient. This allows for the integration of more functionalities in the future, such as 3D reconstruction.
Acknowledgments
We want to thank Laurynas Karazija for evaluating model efficiency, Luke Melas-Kyriazi and Jianyuan Wang for their paper comments, Roman Shapovalov, Iurii Makarov, Shalini Maiti, and Adam W. Harley for the insightful discussions. Christian Rupprecht was supported by ERC-CoG UNION101001212 and VisualAI EP/T028572/1.
Appendix
A Additional ablations
Here we provide additional ablation experiments that supplement the model choices from the main paper.
Support points.
In addition to global and local support grids described in the paper, we also test using the SIFT detector [39] to select support points in Fig. 9. While adding SIFT support points is better than sampling points uniformly on a global regular grid, adding a local grid yields significantly better results. The combination of global and local grids is slightly better, and in Fig. 9 we further test this combination with different grid sizes ranging from 4 $\times$ 4 to 10 $\times$ 10. We find that a configuration of 25 global and 64 local points is optimal for TAP-Vid-DAVIS in terms of speed and accuracy. While there are differences in performance across these hyper-parameters, the overall impact on performance is less than one percentage point for a wide variety of configurations. We use the selected configuration for evaluation on all the TAP-Vid benchmarks.
Figure 8: Accuracy on TAP-Vid-DAVIS depending on the number of additional local, global grid and SIFT support points. When combining point sources, we use (close to) the same number of points.